OpenStudy (anonymous):

Find the real solution of the equation: (5/x+4) = 4 + (3/x-2)

OpenStudy (pdoctor):

Clarify your parentheses

OpenStudy (anonymous):

it's 5 over x+4 and 4 + 3 over x-2

OpenStudy (turingtest):

5/(x+4) you mean?

OpenStudy (pdoctor):

So 5/(x+4) = (4+3)/(x-2)?

OpenStudy (turingtest):

or 4+[3/(x-2)] ?

OpenStudy (anonymous):

turing test you are correct

OpenStudy (anonymous):

the left side of Pdoctor's equation is also correct 5/(x+4)

OpenStudy (turingtest):

ok, multiply both sides by (x+4)(x-2) 5(x-2)=4(x+4)(x-2)+3(x+4) 5x-10=4(x^2+2x-8)+3x+12 2x-22=4x^2+8x-32 4x^2+6x-10=0

OpenStudy (turingtest):

2x^2+3x-5=0 can we factor it or do we have to do the quadratic equation? let me see...

OpenStudy (turingtest):

yeah it factors into (2x+5)(x-1)=0 so x={-5/2,1}

OpenStudy (anonymous):

thank you so much!

OpenStudy (turingtest):

no prob!

OpenStudy (turingtest):

medals are nice though... (hint hint)

OpenStudy (turingtest):

thanks!

OpenStudy (anonymous):

no probelm! wanna help me solve another problem? :D

OpenStudy (turingtest):

let's go

OpenStudy (anonymous):



OpenStudy (turingtest):

absolute value, eh?

OpenStudy (anonymous):

Yeah...do you add and subtract 1/3 to 1?

OpenStudy (turingtest):

\[\left| x/2-1/3 \right|=1\]\[-1=x/2-1/3=1\]\[-2/3=x/2=4/3\]\[-4/3=x=8/3\]\[x=\left\{ -4/3,8/3 \right\}\]

OpenStudy (turingtest):

not until you convert it the way I did to -1=x/2-1/3=1 but that's not really the way they like you to write it. thay usually want you to say x/2-1/3=-1 or x/2-1/3=1 and then solve each seperaetly. you get the same answer thoug, but I did a sort of shorthand.

OpenStudy (anonymous):

ok i see what you mean thank you very much again!

OpenStudy (turingtest):

anytime